OpenStudy (anonymous):

the product of two consecutive integers is 110. find the integers?

OpenStudy (watchmath):

smart guessing? 10, 11 :D

OpenStudy (anonymous):

good guess!

myininaya (myininaya):

x= is first integer x+1=next integer x(x+1)=110

OpenStudy (watchmath):

it's been a while myininaya :D. How are you doing?

OpenStudy (anonymous):

If you want the method to solve them, assume the numbers x and x+1, then: \[x(x+1)=110 \implies x^2+x-110=0 \implies (x-10)(x+11)=0\] So x is either 10 or -11. Hence the numbers are either (10 and 11) or (-11 and -10).

myininaya (myininaya):

x^2+x-110=0 (x-10)(x+11)=0 i'm fine x=10 x=-11 if x=10, then x+1=10+1=11 since (11)(10)=110 you could also do if x=-11, then x+1=-11+1=-10 since (-11)(-10)=110

myininaya (myininaya):

how are you watchmath guy